# The Global Sections of Chiral de Rham Complexes on Closed Complex Curves

### v2

@@ 1. Introduction
The space of global sections $\Gamma(X,\Omega_{X}^{ch})$ is always a vertex superalgebra, and it is known to have extra structure when $X$ is endowed with geometric structures. For example, when $X$ is Kähler or hyperkähler, it has an $\mathcal{N}=2$ structure or $\mathcal{N}=4$ structure, respectively [2]. Recently in [17, 11], Linshaw and the first auther gave a complete description of $\Gamma(X,\Omega_{X}^{ch})$ for any compact Ricci-flat Kähler manifold $X$. In [4], for any congruence subgroup $G\subset SL(2,\mathbb{R})$, Dai constructed a basis of the $G$-invariant global sections of the chiral de Rham complex on the upper half plane, which are holomorphic at the cusps. The vertex operations are determined by a modification of the Rankin-Cohen brackets of modular forms. Further in [5], it is shown that if the global sections are allowed to be meromorphic at cusp, the space of $G$-invariant global sections will be a simple vertex algebra. In this paper, we extend the method developed in [17] to closed local symmetric space and calculate the space of global sections of chiral de Rham complex on any closed curves with genus $g\geq 2$ (Theorem 4.6).
The paper is organized as follows: in secion 2, we introduce the chiral de Rham complex; in section 3, we introduce the method of calculating the global sections of chiral de Rham complex in [17] and apply the method to Hermitian locally symmetric spaces. In section 4, we calculate the space of global sections for any closed complex curve $X$ with genus $g\geq 2$.
@@ 2.1. Vertex algebras
In this paper, we will follow the formalism of vertex algebras developed in [8]. A vertex algebra is the data $({\mathcal{A}},Y,L_{-1},1)$. In this notation,
@@
    $Y$ is an even linear map
    $Y:{\mathcal{A}}\to\operatorname{End}({\mathcal{A}})[[z,z^{-1}]],\qquad Y(a)=a(z)=\sum_{n\in\mathbb{Z}}a_{(n)}z^{-n-1}.$
    Here $z$ is a formal variable and $a(z)$ is called the field corresponding to $a$.
@@
The Wick product of $a(z)$ and $b(z)$ is $:a(z)b(z):\ =(a_{(-1)}b)(z)$. The other negative products are given by
$:\partial^{n}a(z)b(z):\ =n!(a_{(-n-1)}b)(z).$
2.2. $\beta\gamma-bc$ system.
@@
$Q(z)=\sum^{d}_{i=1}:\beta^{i}(z)c^{i}(z):,\quad L(z)=\sum^{d}_{i=1}(:\beta^{i}(z)\partial\gamma^{i}(z):-:b^{i}(z)\partial{c^{i}}(z):),$
@@
Here $\gamma^{i}_{(-1)}$ acts on the analytic function ring $\mathcal{O}(U)$ by function multiplications.
@@
(i) The chiral Hodge cohomology $H^{*}(X,\Omega^{ch}_{X})$ is equal to the cohomology of the complex $(\Omega^{ch,*}_{X},\bar{\partial})$.
@@ 3. Antiholomorphic vector bundles corresponding to chiral de Rham complexes
Let $\tilde{\gamma}^{1},\ldots,\tilde{\gamma}^{d}$ be the new holomorphic local coordinates on $\tilde{U}$ with the relations $\gamma^{i}=g_{i}(\tilde{\gamma})$ and $\tilde{\gamma}^{i}=f_{i}(\gamma)$, the Kähler form is expressed locally as $h|_{\tilde{U}}=\sum_{i,j}\tilde{H}_{ij}d\tilde{\gamma}^{i}\wedge d\tilde{\gamma}^{j}$ in the new coordinate system, where $(\tilde{H}^{jk})$ is the inverse of $(\tilde{H}_{ij})$. Let
@@
So under the change of coordinates, both $\bm{B}^{i}$ and $\bm{b}^{i}$ change like $d\overline{\gamma}^{i}$, and both $\bm{\Gamma}^{i}$ and $\bm{c}^{i}$ change like $\frac{\partial}{\partial\overline{\gamma}^{i}}$.
@@
$S^{\gamma}[k,l,s]=\{A\in S^{\gamma}[k,l]|\text{the number of $\bm{B}$'s in $A$ $-$ the number of $\bm{\varGamma}$'s in $A$ $=$ $s$}\}.$
@@
(3.3) ¯$\bar{\partial}^{\prime}\bm{B}^{i}_{(k)}$ $=\sum_{j}\theta_{ij}\bm{B}^{j}_{(k)},\bar{\partial}^{\prime}\bm{b}^{i}_{(k)}=\sum_{j}\theta_{ij}\bm{b}^{j}_{(k)},$
¯$\bar{\partial}^{\prime}\bm{\Gamma}^{i}_{(k)}$ $=-\sum_{j}\theta_{ji}\bm{\Gamma}^{j}_{(k)},\bar{\partial}^{\prime}\bm{c}^{i}_{(k)}=-\sum_{j}\theta_{ji}\bm{c}^{j}_{(k)}$
@@
Let $\Omega^{0,k}_{X}(SW(\bar{T}^{*}X))$ be the sheaf of smooth (0, k)-forms with values in $SW(\bar{T}^{*}X)$ and let $\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))=\bigoplus\Omega^{0,k}_{X}(SW(\bar{T}^{*}X))$.
@@
Here $F_{i}$, mapping $\Omega^{0,*}_{X}(SW(\bar{T}X)[k,l,m])$ to $\Omega^{0,*+1}_{X}(SW(\bar{T}X)[k,l,m-i])$, is a first differential operator. Locally, for $a\in SW^{\gamma}$, $f\in\Omega^{0,*}_{X}(U)$,
@@
$\varGamma^{i}(z)=\sum_{j\neq 0}\frac{1}{-j}\mathbf{\Gamma}^{i}_{(j)}z^{-j}.$
@@ 3.4. Hermitian locally symmetric space
[…]
@@
$I_{U}(F_{1}(a\otimes f))$ $=$ $\sum_{i,j,k}[::\Gamma^{k}\mathbf{c}^{i}:\mathbf{b}^{j}:_{(0)},a]f\mathbf{B}^{k}_{(0)}\theta_{ij}$
@@
for any $m\in\mathbb{Z}$, and
$\sum_{j}\mathbf{B}^{j}_{(n)}\otimes\frac{\partial}{\partial\bar{\gamma}^{j}},\quad n\neq 0$
From the expression of $F_{1}$ in equation (3.7) and the fact that $\nabla R=0$, we can see that $[\nabla,F_{1}]=0$. ∎
Locally, let $\tilde{F}_{2}^{j}$ be the operator on $\Omega^{0,*}(SW(\bar{T}^{*}X))$ such that
@@
$[\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},F_{2}]=0,\quad[\nabla_{\frac{\partial}{\partial\bar{\gamma}^{i}}},F_{2}]=\tilde{F}_{2}^{j}\tilde{R}(H^{jl}\frac{\partial}{\partial\gamma^{l}},\frac{\partial}{\partial\bar{\gamma}^{i}}).$
@@
By definition of $\tilde{F}_{2}^{j}$ and $\nabla R=0$, $\nabla$ commutes with $\tilde{F}_{2}^{j}\otimes H^{jl}\frac{\partial}{\partial\gamma^{l}}.$ Since
$[\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},\nabla^{1,0}]=0,$
@@
$[\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},F_{2}]=[\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},\sum_{j,l}\tilde{F}_{2}^{j}H^{jl}\nabla_{\frac{\partial}{\partial\gamma^{l}}}]=0.$
$[\nabla_{\frac{\partial}{\partial\bar{\gamma}^{i}}},F_{2}]=[\nabla_{\frac{\partial}{\partial\bar{\gamma}^{i}}},\sum_{j,l}\tilde{F}_{2}^{j}H^{jl}\nabla_{\frac{\partial}{\partial\gamma^{l}}}]=\tilde{F}_{2}^{j}\tilde{R}(H^{jl}\frac{\partial}{\partial\gamma^{l}},\frac{\partial}{\partial\bar{\gamma}^{i}}).$
When $X$ be a closed complex curve with genus $g\geq 2$. We use the upper half plane as a covering space $\tilde{X}$ of $X$, then a Kähler form $h$ on $X$ is locally given by: $h|_{U}=Hd\gamma\wedge d\bar{\gamma}=\frac{i}{2y^{2}}d\gamma\wedge d\bar{\gamma}$, where $y=\operatorname{Im}\gamma>0$ and $\gamma$ is a holomorphic local coordinate. The generators of $\mathcal{W}_{+}^{\gamma}$ are $\mathbf{\Gamma},\mathbf{c},\mathbf{b},\mathbf{B}$. In this case $X$ is a Hermitian locally symmetric space,
@@
Here $F_{2}=\tilde{F}_{2}H^{-1}\nabla_{\frac{\partial}{\partial{\gamma}}}$ and $\tilde{F}_{2}$ commutes with $\nabla_{\frac{\partial}{\partial\gamma}}$ as in the proof of Lemma 3.3. ∎
@@ 4. Global Sections on closed Complex Curves
In this section we use Theorem 3.2 to calculate the space of global sections of chiral de Rham complex on closed curves with genus $g\geq 2$.
4.1. The curvature form on $SW(\overline{T}^{*}X)$.
@@
$-iH^{-1}R(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})\frac{\partial}{\partial\gamma}=-\frac{\partial}{\partial\gamma}.$
So by equation (3.1) we have:
(4.1) $-iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})A=-A,\quad\text{if }A=\bm{B}_{(k)}\text{ or }A=\bm{b}_{(k)}.$
@@
(4.2) $-iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})A=A,\quad\text{ if }A=\bm{\Gamma}_{(k)}\text{ or }A=\bm{c}_{(k)}.$
Then we obtain the following result:
@@
(4.3) $-iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})a_{s}=(l-s)a_{s}$
@@
$=\text{the number of $\bm{B}$'s}-\text{the number of $\bm{\Gamma}$'s}.$ ’s .
@@
where $a_{s-i}$ is a smooth section of $SW(\bar{T}^{*}X)[k,l,s-i]$ and $a_{s}$ is nonzero. The summation is finite since $SW(\bar{T}^{*}X)[k,l,s-i]=\mathbf{0}$ when $k<i-s$. By Corollary 3.1 , $\bar{D}a=0$ is divided into a series of equations bellow:
@@
Case 1: $l-s<0$
The following vanishing theorem for holomorphic sections of Hermitian vector bundles is well known (For a reference, see Chapter III of [10]),
@@
When $l-s<0$, by Lemma 4.1, the mean curvature of $SW(\bar{T}^{*}X)[k,l,s]$ is negative-definite everywhere on $X$, so by above Theorem 4.1, there is no nonzero holomorphic section $a_{s}$.
Case 2: $l-s>0$.
Let $a_{s}$ a holomorphic section of $SW(\bar{T}^{*}X)[k,l,s]$ with $l-s>0$. Let $\bar{\partial}^{\prime*}$ be the dual operator of $\bar{\partial}^{\prime}$, then ¯$\bar{\partial}^{\prime*}=-iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}$.
@@
Given a nonzero holomorphic section $a_{s}$ of $SW(\bar{T}^{*}X)[k,l,s]$ with $l-s>0$, $a_{s-1}=\frac{-1}{l-s}\bar{\partial}^{\prime*}F_{1}a_{s}$ is a solution of the equation ¯$\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$.
@@
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s}$ $=-id\bar{\gamma}H^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{1}a_{s}$
@@
Suppose $a_{s-1}$ is given as in Lemma 4.2, then $a_{s-2}=\frac{-1}{2l-2s+1}(\bar{\partial}^{\prime*}F_{1}a_{s-1}+\bar{\partial}^{\prime*}F_{2}a_{s})$ is a solution of the equation ¯$\bar{\partial}^{\prime}a_{s-2}+F_{1}a_{s-1}+F_{2}a_{s}=0$.
@@
Since ¯$\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$, we have:
@@
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-1}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{1}a_{s-1}$
@@
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-1}=(2l-2s+1)F_{1}a_{s-1}.$
For the term ¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s}$, using Lemma 3.4, we have:
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{2}a_{s}$
@@
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s}$ $=(l-s)F_{2}a_{s}+(l-s+1)F_{2}a_{s}$

### v3

@@ 1. Introduction
The space of global sections $\Gamma(X,\Omega_{X}^{ch})$ is always a vertex superalgebra, and it is known to have extra structure when $X$ is endowed with geometric structures. For example, when $X$ is Kähler or hyperkähler, it has an $\mathcal{N}=2$ structure or $\mathcal{N}=4$ structure, respectively [2]. Recently in [18, 12], Linshaw and the first author gave a complete description of $\Gamma(X,\Omega_{X}^{ch})$ for any compact Ricci-flat Kähler manifold $X$. In [4], for any congruence subgroup $G\subset SL(2,\mathbb{R})$, Dai constructed a basis of the $G$-invariant global sections of the chiral de Rham complex on the upper half plane, which are holomorphic at the cusps. The vertex operations are determined by a modification of the Rankin-Cohen brackets of modular forms. Further in [5], it is shown that if the global sections are allowed to be meromorphic at cusp, the space of $G$-invariant global sections will be a simple vertex algebra. There have been some descriptions about the structure of $H^{0}(X,\Omega^{ch}_{X})$ when $X$ is a closed complex curve with genus 0 or $1$. In the case $g=0$, $X$ is of constant positive curvature and a description of $H^{0}(X,\Omega^{ch}_{X})$ through $\hat{sl_{2}}-$modules was given in [14]. In the case $g=1$, $X$ is of 0 curvature and $H^{0}(X,\Omega^{ch}_{X})$ is just a $\beta\gamma-bc$ system. The case $g\geq 2$ has never been calculated yet, with the curvature of $X$ being constant negative.
When $X$ is a compact Kähler manifold, a filtration $(\mathcal{F}_{s})_{s\in\mathbb{Z}}$ of $H^{0}(X,\Omega^{ch}_{X})$ was in fact constructed in [18], such that $\bigoplus_{s}\mathcal{F}_{s}/\mathcal{F}_{s-1}$ is isomorphic to the space of holomorphic sections of $SW(\bar{T}^{*}X)$, where $SW(\bar{T}^{*}X)$ is isomorphic to
$\operatorname{Sym}^{*}(\bigoplus^{\infty}_{n=1}\bar{T}X)\bigotimes\operatorname{Sym}^{*}(\bigoplus^{\infty}_{n=1}\bar{T}^{*}X)\bigotimes\wedge^{*}(\bigoplus^{\infty}_{n=1}\bar{T}X)\bigotimes\wedge^{*}(\bigoplus^{\infty}_{n=1}\bar{T}^{*}X)$
[…]
The structure of $H^{0}(X,\Omega^{ch}_{X})$ when $X$ is a compact Ricci-flat Kähler manifold was calculated in [[18]]. In this paper, we extend the method developed in [18] to a class of compact non Ricci-flat Kähler manifolds, that is, the closed locally symmetric spaces, and calculate the space of global sections of chiral de Rham complex on any closed curves with genus $g\geq 2$ (Theorem 4.6).
The paper is organized as follows: in section 2, we introduce the chiral de Rham complex; in section 3, we introduce the method of calculating the global sections of chiral de Rham complex in [18] and apply the method to Hermitian locally symmetric spaces. In section 4, we calculate the space of global sections for any closed complex curve $X$ with genus $g\geq 2$.
@@ 2.1. Vertex algebras
In this paper, we will follow the formalism of vertex algebras developed in [9].
A formal distribution is the following formal expressions
$\sum_{m,n,\ldots\in\mathbb{Z}}a_{m,n,\ldots}z^{m}w^{n}\ldots,$
where $a_{m,n,\ldots}$ are elements of a vector space $U$ over $\mathbb{C}$, and $z,w,\ldots$ are indeterminates. These expressions form a vector space over $\mathbb{C}$ denoted by $U[[z,z^{-1},w,w^{-1},\ldots]]$.
A vertex algebra is the data $({\mathcal{A}},Y,L_{-1},1)$. In this notation,
@@
    $Y$ is an even linear map Y : A→End(A) [[z, z^-1]],  Y(a) = a(z) = ∑_n∈Za_(n)z^-n-1. Here $z$ is a formal variable and $a(z)$ is called the field corresponding to $a$.
@@
The Wick product of $a(z)$ and $b(z)$ is $a(z)b(z):\ =(a_{(-1)}b)(z)$. The other negative products are given by
$\partial^{n}a(z)b(z):\ =n!(a_{(-n-1)}b)(z).$
Note that for any formal distributions $a(z)=\sum_{n\in\mathbb{Z}}a_{(n)}z^{-n-1}$ and $b(w)=\sum_{n\in\mathbb{Z}}b_{(n)}w^{-n-1}$ which are not necessarily fields of vertex algebras, the product $a_{(n)}b$, $n\in\mathbb{Z}$ can still be defined.
$2.$ $\beta\gamma-bc$ system.
@@
(2.1) $Q(z)=\sum^{d}_{i=1}:\beta^{i}(z)c^{i}(z):,\quad L(z)=\sum^{d}_{i=1}(:\beta^{i}(z)\partial\gamma^{i}(z):-:b^{i}(z)\partial{c^{i}}(z):),$
@@
Here the action of $\gamma^{i}_{(-1)}$ on the analytic function ring $\mathcal{O}(U)$ is identified with the product of $\gamma^{i}$.
@@
[…]
@@ 3. Antiholomorphic vector bundles corresponding to chiral de Rham complexes
[…]
@@
So under the change of coordinates, both $\bm{B}^{i}$ and $\bm{b}^{i}$ change like $d\overline{\gamma}^{i}$, and both $\bm{\Gamma}^{i}$ and $\bm{c}^{i}$ change like $\frac{\partial}{\partial\bar{\gamma}^{i}}$.
@@
[…]
@@
(3.3) $\bar{\partial}^{\prime}\bm{B}^{i}_{(k)}$ $=\sum_{j}\theta_{ij}\bm{B}^{j}_{(k)},\bar{\partial}^{\prime}\bm{b}^{i}_{(k)}=\sum_{j}\theta_{ij}\bm{b}^{j}_{(k)},$
$\bar{\partial}^{\prime}\bm{\Gamma}^{i}_{(k)}$ $=-\sum_{j}\theta_{ji}\bm{\Gamma}^{j}_{(k)},\bar{\partial}^{\prime}\bm{c}^{i}_{(k)}=-\sum_{j}\theta_{ji}\bm{c}^{j}_{(k)}$
@@
Let $\Omega^{0,k}_{X}$ be the sheaf of smooth $(0,k)$-forms with values in $\mathbb{C}$ and let $\Omega^{0,*}_{X}=\bigoplus\Omega^{0,k}_{X}$. Let $\Omega^{0,k}_{X}(SW(\bar{T}^{*}X))$ be the sheaf of smooth (0, k)-forms with values in $SW(\bar{T}^{*}X)$ and let $\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))=\bigoplus\Omega^{0,k}_{X}(SW(\bar{T}^{*}X))$.
@@
Here $F_{i}$, mapping $\Omega^{0,*}_{X}(SW(\bar{T}^{*}X)[k,l,m])$ to $\Omega^{0,*+1}_{X}(SW(\bar{T}^{*}X)[k,l,m-i])$, is a first order differential operator. Locally, for $a\in SW^{\gamma}$, $f\in\Omega^{0,*}_{X}(U)$,
@@
$\varGamma^{i}(z)=\sum_{j\neq 0}\frac{1}{-j}\mathbf{\Gamma}^{i}_{(j)}z^{-j},$
which is a formal distribution, but is not a field in a vertex algebra for any $i$.
@@ 3.4. Hermitian locally symmetric space
Here $R(\cdot,\cdot):TX\to TX$ is the curvature (1,1)-form, and $\langle\cdot,\cdot\rangle$ is the contraction of tangent field with $1$-form.
@@
$I_{U}(F_{1}(a\otimes f))$ $=$ $\sum_{i,j,k}[::\varGamma^{k}\mathbf{c}^{i}:\mathbf{b}^{j}:_{(0)},a]f\mathbf{B}^{k}_{(0)}\theta_{ij}$
@@
for any $m\in\mathbb{Z},$ and commutes with
$\sum_{j}\mathbf{B}^{j}_{(n)}\otimes\frac{\partial}{\partial\bar{\gamma}^{j}},\quad n\neq 0.$
From the expression of $F_{1}$ in equation (3.7) and the fact that $\nabla R=0$, we can see that $\nabla,F_{1}]=0$. ∎
For the convenience of deriving some useful formulas about $F_{2}$, we introduce an operator $\tilde{F}_{2}^{j}$ on $\Omega^{0,*}(SW(\bar{T}^{*}X))$ which is locally given by:
@@
$\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},F_{2}]=0,\quad[\nabla_{\frac{\partial}{\partial\bar{\gamma}^{i}}},F_{2}]=\tilde{F}_{2}^{j}\tilde{R}(H^{jl}\frac{\partial}{\partial\gamma^{l}},\frac{\partial}{\partial\bar{\gamma}^{i}}).$
@@
By the definition of $\tilde{F}_{2}^{j}$ and the condition $\nabla R=0$, $\nabla$ commutes with $\tilde{F}_{2}^{j}\otimes H^{jl}\frac{\partial}{\partial\gamma^{l}}.$ Since
$\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},\nabla^{1,0}]=0,$
@@
$\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},F_{2}]=[\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},\sum_{j,l}\tilde{F}_{2}^{j}H^{jl}\nabla_{\frac{\partial}{\partial\gamma^{l}}}]=0.$
$\nabla_{\frac{\partial}{\partial\bar{\gamma}^{i}}},F_{2}]=[\nabla_{\frac{\partial}{\partial\bar{\gamma}^{i}}},\sum_{j,l}\tilde{F}_{2}^{j}H^{jl}\nabla_{\frac{\partial}{\partial\gamma^{l}}}]=\tilde{F}_{2}^{j}\tilde{R}(H^{jl}\frac{\partial}{\partial\gamma^{l}},\frac{\partial}{\partial\bar{\gamma}^{i}}).$
When $X$ is a closed complex curve with genus $g\geq 2$, it can be constructed as a quotient $\mathbb{H}/\Gamma$, here $\mathbb{H}=\left\{z\in\mathbb{C}|\operatorname{Im}(z)>0\right\}$ and $\Gamma$ is a discrete subgroup of the holomorphic automorphism group $\operatorname{PSL}(2,\mathbb{R})=\left\{z\mapsto\frac{az+b}{cz+d}|a,b,c,d\in\mathbb{R},ad-bc=1\right\}$. Moreover, $\Gamma$ is required to act properly discontinuously without fixed points on $\mathbb{H}$ to make sure that $\mathbb{H}/\Gamma$ is compact. We use the upper half plane $\mathbb{H}$ as a covering space $\tilde{X}$ of $X$, and a Kähler form $h$ on $X$ is locally given by: $h|_{U}=Hd\gamma\wedge d\bar{\gamma}=\frac{i}{2y^{2}}d\gamma\wedge d\bar{\gamma}$, where $y=\operatorname{Im}\gamma>0$ and $\gamma$ is a holomorphic local coordinate. The generators of $\mathcal{W}_{+}^{\gamma}$ are $\mathbf{\Gamma},\mathbf{c},\mathbf{b},\mathbf{B}$. In this case $X$ is a Hermitian locally symmetric space,
@@
Here $F_{2}=\tilde{F}_{2}H^{-1}\nabla_{\frac{\partial}{\partial{\gamma}}}$ and $\tilde{F}_{2}$ commutes with $\nabla$ as in the proof of Lemma 3.3. ∎
@@ 4. Global Sections on closed Complex Curves
In this section we use Theorem 3.2 to calculate the space of global sections of chiral de Rham complex on closed complex curves with genus $g\geq 2$. Let $X$ be a closed complex curves with genus $g\geq 2$. As is mentioned in section 3.4, $X$ is uniformized by $\mathbb{H}$ and we choose the canonical metric induced by the Kähler form $h$ : $h|_{U}=Hd\gamma\wedge d\bar{\gamma}=\frac{i}{2y^{2}}d\gamma\wedge d\bar{\gamma}$, where $y=\operatorname{Im}\gamma>0$ and $\gamma$ is a holomorphic local coordinate. The curvature of $X$ corresponding to this canonical metric is constant negative.
4.1. The curvature form on $SW(\bar{T}^{*}X)$.
@@
$iH^{-1}R(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})\frac{\partial}{\partial\gamma}=-\frac{\partial}{\partial\gamma}.$
We have:
(4.1) $iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})A=-A,\quad\text{if }A=\bm{B}_{(k)}\text{ or }A=\bm{b}_{(k)},$
which follows from the fact that the antiholomorphic cotangent bundle $\bar{T}^{*}X$ of $X$ has curvature 1.-1.
@@
(4.2) $iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})A=A,\quad\text{ if }A=\bm{\Gamma}_{(k)}\text{ or }A=\bm{c}_{(k)},$
which follows from the fact that the antiholomorphic tangent bundle $\bar{T}X$ of $X$ has curvature $1.$
We obtain the following result:
@@
(4.3) $iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})a_{s}=(l-s)a_{s}.$
Therefore, the curvature of the sub bundle $SW(\bar{T}^{*}(X))[k,l,s]$ of $SW(\bar{T}^{*}(X))$ is just lsl-s.
@@
$s=\text{the number of $\bm{B}$'s}-\text{the number of $\bm{\Gamma}$'s}.$
@@
where $a_{s-i}$ is a smooth section of $SW(\bar{T}^{*}X)[k,l,s-i]$ and $a_{s}$ is nonzero. The summation is finite since $SW(\bar{T}^{*}X)[k,l,s-i]=\mathbf{0}$ when k<isk<i-s. By Corollary 3.1 , $\bar{D}a=0$ is divided into a series of equations bellow:
@@
Case 1: ls<0l-s<0
The following vanishing theorem for holomorphic sections of Hermitian vector bundles is well known (cf. [ [11], III, 1.9(ii)]),
@@
When ls<0l-s<0, by Lemma 4.1, the mean curvature of $SW(\bar{T}^{*}X)[k,l,s]$ is negative-definite everywhere on $X$, so by above Theorem 4.1, there is no nonzero holomorphic section $a_{s}$.
Case 2: ls>0l-s>0.
Let $a_{s}$ be a holomorphic section of $SW(\bar{T}^{*}X)[k,l,s]$ with ls>0l-s>0. Let $\bar{\partial}^{\prime*}$ be the dual operator of $\bar{\partial}^{\prime}$, then $\bar{\partial}^{\prime*}=-iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}$.
@@
Given a nonzero holomorphic section $a_{s}$ of $SW(\bar{T}^{*}X)[k,l,s]$ with ls>0l-s>0, $a_{s-1}=\frac{-1}{l-s}\bar{\partial}^{\prime*}F_{1}a_{s}$ is a solution of the equation $\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$.
@@
$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s}$ $=-id\bar{\gamma}H^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{1}a_{s}$
@@
Suppose $a_{s-1}$ is given as in Lemma 4.2, then $a_{s-2}=\frac{-1}{2l-2s+1}(\bar{\partial}^{\prime*}F_{1}a_{s-1}+\bar{\partial}^{\prime*}F_{2}a_{s})$ is a solution of the equation $\bar{\partial}^{\prime}a_{s-2}+F_{1}a_{s-1}+F_{2}a_{s}=0$.
@@
Since $\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$, we have:
@@
$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-1}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{1}a_{s-1}$
@@
$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-1}=(2l-2s+1)F_{1}a_{s-1}.$
For the term $\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s}$, using Lemma 3.4, we have:
$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{2}a_{s}$
@@
$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s}$ $=(l-s)F_{2}a_{s}+(l-s+1)F_{2}a_{s}$